# a data-driven parametrization of the CMB temperature power spectrum

### v1

@@ I Introduction
A related approach was pursued in Ref. [15], where it was shown that measurements of the CMB power spectrum can be understood in terms of a phenomenological representation of four variables. Our approach is similar, except that our reparametrization of the CMB features (and therefore of the underlying cosmological parameters) is performed by a neural network and is thus entirely data-driven, as well as non-linear. Refs. [16, 17, 18, 19, 20] also investigated a data-driven parametrization of cosmological models based on principal component analysis (PCA), a linear compression scheme, while Ref. [21] explored the use of a VAE to compress $w$CDM cosmologies using matter power spectra.
@@ III.2 Variational autoencoders
VAEs are unsupervised encoder-decoder networks that learn to compress the input data into a lower-dimensional representation, known as the latent representation or latent variables, and then use these latents to reconstruct something that is closely similar to the input data [49, 13, 14]. The former part of the algorithm is the encoder and the latter the decoder; we choose the network architectures of the encoder and decoder to be simple 1D convolutional neural networks. The latent representation aims to capture all the relevant information required to reconstruct the input. The latent representation of a given input $\bm{x}$ is a probability distribution function $p(\bm{z}|\bm{x})$ which is usually represented by a multivariate diagonal Gaussian $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}$ and $\bm{\sigma}$ are the means and standard deviations of the Gaussian distribution of each latent parameter $z$. The size of the vectors $\bm{\mu}$ and $\bm{\sigma}$ is $L$, namely the latent space dimensionality. The means and standard deviations of each latent dimension are the outputs of the encoder, while the decoder takes as input samples $\bm{z}\sim\mathcal{N}(\bm{\mu},\bm{\sigma})$, thus returning a distribution of reconstructed outputs $\bm{\hat{x}}$ from a single input $\bm{x}$.
@@ IV.2.2 Latent parameter constraints from Planck data
We compare the latent posterior constraints against several theoretical expectations. In both the right and left panels, we show the range of latent values corresponding to a single cosmology – a $\Lambda$CDM cosmology with cosmological parameters set by the Planck best-fit values – in green. To obtain the green latent distribution, we take the best-fit $\Lambda$CDM cosmological parameters from Ref. [2], use CLASS to generate $D_{\ell}^{\rm TT}$ and use the trained VAE encoder to map $D_{\ell}^{\rm TT}$ to its encoded $L$-dimensional Gaussian distribution. The comparison shows that our latent posterior constraints are consistent with the latent values corresponding to the best-fit $\Lambda$CDM cosmology from Planck obtained from a traditional Bayesian approach. This is the case for both the $\Lambda$CDM latents and the EDE latents.
In the right panel, we additionally compare our constraints to the latent values corresponding to the best-fit EDE cosmology under Plik_lite TT data (including a Planck-informed prior on $\tau=0.0506\pm 0.0086$ [2]). We obtain the best-fit EDE cosmology by running a global minimization with the simulated-annealing minimizer pinc [56], yielding the best-fit values $f_{\mathrm{EDE}}=0.06$, $\log z_{\mathrm{c}}=3.4$, $\theta_{\mathrm{i}}=2.4$. We find that the latent posterior constraints are consistent with both the best-fit $\Lambda$CDM cosmology reported by Ref. [2] and the best-fit EDE cosmology under Plik_lite. Our results are therefore consistent with previous constraints obtained with traditional parameter inference techniques using similar CMB data. [^4] Our constraints are based on the Plik_lite TT likelihood, while previous constraints were based on the full Plik likelihood (combined with other data). We used the Plik_lite TT likelihood because it allows for $D_{\ell}^{\mathrm{TT}}$ as input rather than requiring cosmological and nuisance parameters. We verified that Plik_lite TT gives comparable (albeit slightly looser constraints) on EDE than Plik TT using MontePython [57, 58].
Although the two cosmologies are mildly separated in latent space, the CMB temperature power spectrum alone is unable to differentiate between those two models, yielding constraints that are consistent with both theoretical expectations. This is not surprising since EDE was constructed in such a way as to preserve the fit to CMB data while allowing for higher values of $H_{0}$. In order to probe the ability of EDE to resolve the Hubble tension in latent space, an inclusion of direct measurements of $H_{0}$ (e.g. [59, 60, 61, 62, 63]) into the training process is necessary, which is left to future work.
@@
V Cosmological information
in latent space
@@ V.1 Mutual information
We calculate MI using the GMM-MI package [65], [^5] https://github.com/dpiras/GMM-MI which fits a Gaussian mixture model to the joint distribution of $x$ and $y$ samples to provide a robust estimate of MI along with its associated uncertainty via bootstrapping. Previous work has already demonstrated the utility of MI in the physical interpretation of latent spaces in the context of predicting the properties of final cosmic structures such as (sub)halo density profiles [66, 67, 68] and the halo mass function [69]. We also use MI to assess the disentanglement of the latent variables in tuning $\beta$ for each VAE: we find that the maximum value of MI between pairs of latents is $\mathcal{O}(10^{-2})$ nat, significantly smaller than the MI between latents and cosmological parameters, thus confirming the disentanglement.
@@ VI Conclusions
Our method is broadly generalizable, enabling us to identify which parameters the data is sensitive to through a data-driven, non-linear approach. Focusing on the well-established cosmological probe of the CMB TT power spectrum allowed us to validate our model and explore its capabilities in a controlled environment. We specifically focused on EDE as it is an example of a phenomenological description of a beyond-standard model of cosmology, which poses challenges related to prior volume effects when performing standard Bayesian analyses. Our methodology also holds promise for compressing other cosmological probes, particularly those related to the late-time Universe, which typically rely on large numbers of correlated parameters. These include, for example, the galaxy power spectrum under the effective field theory of large-scale structure (EFTofLSS, e.g. [75, 76, 77, 78]), which involves many nuisance parameters that can impact the constraints [79, 80, 81, 82]. In future work, we plan to incorporate additional data vectors including CMB polarization and late-time probes to further evaluate the benefits of our approach.
@@
We will make data and materials supporting the results presented in this paper available upon reasonable request.
@@ Acknowledgments
LH thanks Graeme Addison and Charles Bennett for helpful discussions. LLS thanks Elisa Ferreira for insightful discussions. DP was supported by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”. LH was supported a William H. Miller fellowship. EK was supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy: Grant No. EXC-2094 - 390783311. Some computations underlying this work were performed on the Baobab cluster at the University of Geneva, while other parts of this work were performed on the freya cluster maintained by the Max Planck Computing & Data Facility.
@@ Appendix A Comparison between latent posteriors and priors
Fig. 11 compares the latent posterior constraints with the prior volume of the latent space shown as gray contours. The prior contours were generated by encoding all the test set spectra into their respective 5-dimensional (8-dimensional) Gaussian latent distribution predicted by the VAEΛCDM (VAEEDE) encoder, and sampling from each of those multivariate Gaussians once in order to construct the gray contours. This shows that the latent space is very well constrained compared to the range of possible values of the test set cosmologies. The only exceptions are latent 4, 6, 7, 8 for the EDE case (right panel), which carry very little cosmological information about the spectra (as we demonstrate in Sec. V). As a result, the blue marginalized distributions for those latents are close to their respective prior distributions.

### v2

@@ I Introduction
A related approach was pursued in Ref. [15], where it was shown that measurements of the CMB power spectrum can be understood in terms of a phenomenological representation of four variables. Our approach is similar, except that our reparametrization of the CMB features (and therefore of the underlying cosmological parameters) is performed by a neural network and is thus entirely data-driven, as well as non-linear. Refs. [16, 17] additionally explored physical parameter combinations that break degeneracies and accelerate the computation of CMB spectra, Refs. [18, 19, 20, 21, 22] investigated data-driven parametrizations of cosmological models based on principal component analysis (PCA), a linear compression scheme, while Ref. [23] used of a VAE to compress $w$CDM cosmologies using matter power spectra.
@@ III.2 Variational autoencoders
VAEs are unsupervised encoder-decoder networks that learn to compress the input data into a lower-dimensional representation, known as the latent representation or latent variables, and then use these latents to reconstruct something that is closely similar to the input data [51, 13, 14]. The former part of the algorithm is the encoder and the latter the decoder; we choose the network architectures of the encoder and decoder to be simple 1D convolutional neural networks. The latent representation aims to capture all the relevant information required to reconstruct the input. The latent representation of a given input $\bm{x}$ is a probability distribution function $p(\bm{z}|\bm{x})$ which is usually represented by a multivariate diagonal Gaussian $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}$ and $\bm{\sigma}$ are the means and standard deviations of the Gaussian distribution of each latent parameter $z$. The size of the vectors $\bm{\mu}$ and $\bm{\sigma}$ is $L$, namely the latent space dimensionality. The means and standard deviations of each latent dimension are the outputs of the encoder, while the decoder takes as input samples $\bm{z}\sim\mathcal{N}(\bm{\mu},\bm{\sigma})$, thus returning a distribution of reconstructed outputs $\bm{\hat{x}}$ from a single input $\bm{x}$. A VAE can be thought of as a generalization of PCA to non-linear compressions: while PCA performs a linear decomposition of the original input space, the VAE performs a non-linear compression via the non-linear convolutional layers of the encoder.
@@ IV.2.2 Latent parameter constraints from Planck data
We compare the latent posterior constraints against several theoretical expectations. By ‘theoretical expectation’, we mean the range of latent values corresponding to a single cosmology i.e. a single set of cosmological parameter values. The encoder always returns a distribution in latent space, so even for a single set of cosmological parameters one obtains a distribution in latent space (and hence the 2D contours in Fig. 5). In the right and left panels of Fig. 5, we show in green the range of latent values corresponding to a single cosmology (the ‘theoretical expectations’), that is a $\Lambda$CDM cosmology with cosmological parameters set by the Planck best-fit values. To obtain the green latent distribution, we take the best-fit $\Lambda$CDM cosmological parameters from Ref. [2], use CLASS to generate $D_{\ell}^{\rm TT}$ and use the trained VAE encoder to map $D_{\ell}^{\rm TT}$ to its encoded $L$-dimensional Gaussian distribution. The comparison shows that our latent posterior constraints are consistent with the latent values corresponding to the best-fit $\Lambda$CDM cosmology from Planck obtained from a traditional Bayesian approach. This is the case for both the $\Lambda$CDM latents and the EDE latents.
In the right panel, we additionally compare our constraints to another theoretical expectation value; again, this is given by the latent values corresponding to a single cosmology. This time, we consider the latent values corresponding to the best-fit EDE cosmology under Plik_lite TT data (including a Planck-informed prior on $\tau=0.0506\pm 0.0086$ [2]). These are shown as pink contours in the right panel of Fig. 5. We obtain the best-fit EDE cosmology by running a global minimization with the simulated-annealing minimizer pinc [58], yielding the best-fit values $f_{\mathrm{EDE}}=0.06$, $\log z_{\mathrm{c}}=3.4$, $\theta_{\mathrm{i}}=2.4$. We find that the latent posterior constraints are consistent with both the best-fit $\Lambda$CDM cosmology reported by Ref. [2] and the best-fit EDE cosmology under Plik_lite. Our results are therefore consistent with previous constraints obtained with traditional parameter inference techniques using similar CMB data. [^4] Our constraints are based on the Plik_lite TT likelihood, while previous constraints were based on the full Plik likelihood (combined with other data). We used the Plik_lite TT likelihood because it allows for $D_{\ell}^{\mathrm{TT}}$ as input rather than requiring cosmological and nuisance parameters. We verified that Plik_lite TT gives comparable (albeit slightly looser constraints) on EDE than Plik TT using MontePython [59, 60].
Although the two cosmologies (denoted as ‘theoretical expectations’) are mildly separated in latent space, the CMB temperature power spectrum alone is unable to differentiate between those two models, yielding constraints that are consistent with both cosmologies. This is not surprising since EDE was constructed in such a way as to preserve the fit to CMB data while allowing for higher values of $H_{0}$. In order to probe the ability of EDE to resolve the Hubble tension in latent space, an inclusion of direct measurements of $H_{0}$ (e.g. [61, 62, 63, 64, 65]) into the training process is necessary, which is left to future work.
@@
V Cosmological information in latent space
@@ V.1 Mutual information
We calculate MI using the GMM-MI package [67], [^5] https://github.com/dpiras/GMM-MI which fits a Gaussian mixture model to the joint distribution of $x$ and $y$ samples to provide a robust estimate of MI along with its associated uncertainty via bootstrapping. Previous work has already demonstrated the utility of MI in the physical interpretation of latent spaces in the context of predicting the properties of final cosmic structures such as (sub)halo density profiles [68, 69, 70] and the halo mass function [71].
As we show in Sec. V.2 and Sec. V.3, we use MI to quantify the shared information between each latent parameter and each cosmological parameter; samples from their joint distribution are given by the set of cosmological parameters and latents corresponding to the same spectra. We also use MI to assess the disentanglement of the latent variables in tuning $\beta$ for each VAE: we find that the maximum value of MI between pairs of latents is $\mathcal{O}(10^{-2})$ nat, significantly smaller than the MI between latents and cosmological parameters, thus confirming the disentanglement.
@@ VI Conclusions
Our method is broadly generalizable, enabling us to identify which parameters the data is sensitive to through a data-driven, non-linear approach. Focusing on the well-established cosmological probe of the CMB TT power spectrum allowed us to validate our model and explore its capabilities in a controlled environment. We specifically focused on EDE as it is an example of a phenomenological description of a beyond-standard model of cosmology, which poses challenges related to prior volume effects when performing standard Bayesian analyses. Our methodology also holds promise for compressing other cosmological probes, particularly those related to the late-time Universe, which typically rely on large numbers of correlated parameters. These include, for example, the galaxy power spectrum under the effective field theory of large-scale structure (EFTofLSS, e.g. [77, 78, 79, 80]), which involves many nuisance parameters that can impact the constraints [81, 82, 83, 84]. In future work we will investigate whether our method can provide new insights into current cosmological tensions between different datasets by revealing which features of the data are responsible for such discrepancies. This will include the incorporation of additional data vectors, including CMB polarization and late-time probes, to further evaluate the benefits of our approach.
@@
We will make data and materials supporting the results presented in this paper available upon reasonable request. The trained models and an example notebook are available at https://github.com/dpiras/VAExEDE.
@@ Acknowledgments
LH thanks Graeme Addison and Charles Bennett for helpful discussions. LLS thanks Elisa Ferreira for insightful discussions. DP was supported by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”. LH was supported a William H. Miller fellowship. LLS acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 „Quantum Universe“ – 390833306. EK was supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy: Grant No. EXC-2094 - 390783311. Some computations underlying this work were performed on the Baobab cluster at the University of Geneva, while other parts of this work were performed on the freya cluster maintained by the Max Planck Computing & Data Facility.
@@ Appendix A Comparison between latent posteriors and priors
Fig. 11 compares the latent posterior constraints with the prior volume of the latent space shown as gray contours. The prior contours were generated by encoding all the test set spectra into their respective 5-dimensional (8-dimensional) Gaussian latent distribution predicted by the VAEΛCDM (VAEEDE) encoder, and sampling from each of those multivariate Gaussians once in order to construct the gray contours. This shows that the latent space is very well constrained compared to the range of possible values of the test set cosmologies. The only exceptions are latent 4, 6, 7, 8 for the EDE case (right panel), which carry very little cosmological information about the spectra (as we demonstrate in Sec. V). As a result, the blue marginalized distributions for those latents are close to their respective prior distributions. In particular, we note that the posterior distribution of latent $z_{7}$ in the right panel of Fig. 11 even extends beyond the gray distribution. This is because we use broader uniform priors to infer the posterior constraints, and thus there is additional allowed parameter space than what is shown in gray. This behaviour is a result of the poor constraining power of the data for this subdominant latent.